OpenStudy (anonymous):

The table below shows the values of f(x) and g(x) for different values of x. One of the functions is a quadratic function, and the other is an exponential function. Which function is most likely increasing exponentially? x f(x) g(x) 1 2 2 2 5 4 3 10 8 4 17 16 5 26 32 a) f(x), because it grows slower than g(x) b) f(x), because it eventually intersects g(x) c) g(x), because it grows faster than f(x) d) g(x), because it will not intersect f(x)

OpenStudy (anonymous):

@jim_thompson5910 @campbell_st

OpenStudy (anonymous):

@netlopes1

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

@ShadowLegendX

OpenStudy (anonymous):

please help i have no idea how to do this:((

OpenStudy (netlopes1):

Only the answer? "C"

OpenStudy (campbell_st):

looking at f(x) it looks like \[f(x) = x^2 + 1\] and the other is \[g(x) = 2^x\] which is the exponential... and why... what is happening as x gets larger